# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ G.3 Mirror descent on Gaussians
As the Mirror Descent scheme cannot be computed in closed-form for Bregman potential which are not potential energies, and thus are computationally costly, we propose here to restrain ourselves to the Gaussian setting.
We choose as target distribution $\nu=\mathcal{N}(0,\Sigma)$ for $\Sigma$ a symmetric positive definite matrix in $\mathbb{R}^{10\times 10}$, and thus the functional to be optimized is $\mathcal{F}(\mu)=\int V\mathrm{d}\mu+\mathcal{H}(\mu)$ with $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$. The initial distribution is always chosen as $\mu_{0}=\mathcal{N}(0,I_{d})$. In all cases, the step size is chosen as $\tau=0.01$, and we run the scheme for 1500 iterations. For the target distributions, we sample 20 random covariances of the form $\Sigma=UDU^{T}$ with $D$ evenly spaced in log scale between $1$ and $100$, and $U\in\mathbb{R}^{10\times 10}$ chosen as a uniformly random orthogonal matrix, as in [40], and we report the averaged KL divergence over iterations in Figure 2. We add on Figure 6 the same experiments with targets of the form $\mathcal{N}(0,D)$ where $D$ is a diagonal matrix on $\mathbb{R}^{10\times 10}$ sampled uniformly over $[0,50]^{10}$. We compare here the Forward-Backward (FB) scheme of [40], the ideally preconditioned Forward-Backward scheme (PFB), which uses the closed-form (115) derived in Appendix F with $\Lambda=\Sigma$, the Mirror Descent with negative entropy Bregman potential (NEM), whose closed-form was derived in Section D.3
@@
We also experiment with the KL divergence as Bregman potential (KLM) and the ideally preconditioned KL divergence (PKLM). We observe that, even though the objective is convex relative to the Bregman potential, this scheme does not always converge. It might be due to its gradient which might not always be invertible.color=violet!30,size=] PCA: à rediscuter, parce que c’est aussi possible que ce soit un pbm de schéma du type ss-gradient
@@
We note that using as Bregman potential $\phi_{\mu}(\mathrm{T})=\int\psi\circ\mathrm{T}\mathrm{d}\mu$ for $\psi(x)=\frac{1}{2}x^{T}\Lambda^{-1}x$ is equivalent with using a preconditioner with $h^{*}(x)=\frac{1}{2}x^{T}\Lambda x$.
@@
It is well-known that along the Wasserstein gradient flow of the KL divergence starting from a Gaussian and with a Gaussian target (Ornstein-Uhlenbeck process), the measures stay Gaussian [118]. Thus, the Forward-Backward scheme has Gaussian iterates at each step [101, 40]. We also use a linearly preconditioned Forward-Backward scheme, which closed-form is derived in (115) (Appendix F). For the Bregman potential, we choose $\phi_{\mu}(\mathrm{T})=\int\psi\circ\mathrm{T}\ \mathrm{d}\mu$ for $\psi(x)=\frac{1}{2}x^{T}\Sigma x$. In this situation, $\mathcal{G}(\mu)=\int V\mathrm{d}\mu$ is 1-smooth and 1-convex relative to $\phi_{\mu}$. Thus, we can apply Proposition 23. We refer to Appendix F for more details on the convexity of $\mathcal{H}$.
@@ G.4 Single-cell experiments
We show the benefits of using the polynomial preconditioner over the single-cell datasets for different metrics. The first one considered is the Sliced-Wasserstein distance [95, 16], defined as
@@
The second one considered is the Sinkhorn divergence [47] defined as
@@
$\mathrm{OT}_{\varepsilon}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\ \int\|x-y\|_{2}^{2}\ \mathrm{d}\gamma(x,y)+\varepsilon\mathrm{KL}(\gamma||\mu\otimes\nu).$ (128)
The Wasserstein gradient of $\mathrm{S}_{\varepsilon,2}^{2}$ is simply obtained as the potential [47]. Finally, we also consider the energy distance, defined as
@@ H.1 Proof of Proposition 1
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T}_{\phi_{\mu}}^{\mu,\nu},\mathrm{S})$ $=\phi\big{(}(\mathrm{T}_{\phi_{\mu}}^{\mu,\nu})_{\#}\mu\big{)}-\phi(\mathrm{S}_{\#}\nu)-\int\langle\nabla_{\mathrm{W}_{2}}\phi(\mathrm{S}_{\#}\mu)(y),x-y\rangle\ \mathrm{d}(\mathrm{T}_{\phi_{\mu}}^{\mu,\nu},\mathrm{S})_{\#}\mu(x,y)$ (131)
@@
$\geq\mathrm{W}_{\phi}(\nu,\rho)=\mathrm{d}_{\phi_{\rho}}(\mathrm{U}_{\phi_{\rho}}^{\rho,\nu},\mathrm{Id}).$
In the last line, we used Proposition 13, i.e. that the optimal coupling is of the form $(\mathrm{U}_{\phi_{\rho}}^{\rho,\nu},\mathrm{Id})_{\#}\rho$.
H.2 Proof of Proposition 2
Let $\mathrm{T}_{k+1}=\operatorname*{argmin}_{\mathrm{T}\in L^{2}(\mu_{k})}\ \tau\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{Id})$. Applying the 3-point inequality (Lemma 27) with $\psi(\mathrm{T})=\tau\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}$ which is convex, $\mathrm{T}_{0}=\mathrm{Id}$ and $\mathrm{T}^{*}=\mathrm{T}_{k+1}$, we get for all $\mathrm{T}\in L^{2}(\mu_{k})$,
@@ H.2 Proof of Proposition 2
$\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})=\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})-\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{Id})-\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}_{k+1}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}\leq\beta\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})$ (134)
~$\iff\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})\leq\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{Id})+\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}_{k+1}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\beta\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id}).$
@@
~$\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})$$\leq\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{Id})+\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{Id})-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{T}_{k+1})$ (136)
@@
H.3 Proof of Proposition 3
@@
$\mathrm{W}_{\phi}(\mu^{*},\mu_{k+1})\leq\left(1-\alpha\tau\right)\mathrm{W}_{\phi}(\mu^{*},\mu_{k})\leq\left(1-\alpha\tau\right)^{k+1}\mathrm{W}_{\phi}(\mu^{*},\mu_{0}).$ (143)
H.4 Proof of Proposition 4
@@
$\phi_{\mu_{k+1}}^{h^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\big{)}$ $=\phi_{\mu_{k}}^{h^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}$ (144)
@@
$\quad+\frac{1}{\tau}\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T}_{k+1}),$
@@
Plugging this into (144), we get
@@
H.5 Proof of Proposition 5
Let $\mu^{*}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be the minimizer of $\mathcal{F}$, $k\geq 0$ and $\mathrm{T}=\operatorname*{argmin}_{\mathrm{T}\in L^{2}(\mu_{k}),\mathrm{T}_{\#}\mu_{k}=\mu^{*}}\ \mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T})$. First, observe that since $\mu^{*}$ is the minimizer of $\mathcal{F}$, then $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu^{*})=0$ (see e.g. [69, Theorem 3.1]), and thus $\phi_{\mu_{k}}^{h^{*}}(0)={h^{*}}(0)$. Moreover, it induces that $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T})=\mathcal{F}(\mu_{k})-\mathcal{F}(\mu^{*})$ and $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{T})=\mathcal{F}(\mu_{k+1})-\mathcal{F}(\mu^{*})$.
@@ H.5 Proof of Proposition 5
$\mathcal{F}(\mu_{k+1})-\mathcal{F}(\mu^{*})$ $=\left(1-\tau\alpha\right)\big{(}\mathcal{F}(\mu_{k})-\mathcal{F}(\mu^{*})\big{)}$ (153)
$\leq\left(1-\tau\alpha\right)^{k+1}\big{(}\mathcal{F}(\mu_{0})-\mathcal{F}(\mu^{*})\big{)}.$
H.6 Proof of Lemma 9

### v2

@@ G.3 Mirror descent on Gaussians
As the mirror descent scheme cannot be computed in closed-form for Bregman potentials which are not potential energies, and thus are computationally costly, we propose here to restrain ourselves to the Gaussian setting.
We choose as target distribution $\nu=\mathcal{N}(0,\Sigma)$ for $\Sigma$ a symmetric positive definite matrix in $\mathbb{R}^{10\times 10}$, and the functional to be optimized is $\mathcal{F}(\mu)=\int V\mathrm{d}\mu+\mathcal{H}(\mu)$ with $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$. The initial distribution is always chosen as $\mu_{0}=\mathcal{N}(0,I_{d})$. In all cases, the step size is chosen as $\tau=0.01$, and we run the scheme for 1500 iterations. For the target distributions, we sample 20 random covariances of the form $\Sigma=UDU^{T}$ with $D$ evenly spaced in log scale between $1$ and $100$, and $U\in\mathbb{R}^{10\times 10}$ chosen as a uniformly random orthogonal matrix, as in [43], and we report the averaged KL divergence over iterations in Figure 2. We add on Figure 5 the same experiments with targets of the form $\mathcal{N}(0,D)$ where $D$ is a diagonal matrix on $\mathbb{R}^{10\times 10}$ sampled uniformly over $[0,50]^{10}$. We compare here the Forward-Backward (FB) scheme of [43], the ideally preconditioned Forward-Backward scheme (PFB), which uses the closed-form (116) derived in Appendix F with $\Lambda=\Sigma$, and the Mirror Descent with negative entropy Bregman potential (NEM), whose closed-form was derived in Section D.3, and which we recall:
@@
We also experiment with the KL divergence as Bregman potential (KLM) and the ideally preconditioned KL divergence (PKLM). We observe that, even though the objective is convex relative to the Bregman potential, this scheme does not always converge. It might be due to its gradient which might not always be invertible. We leave further investigations for future works.
@@
We note that using as Bregman potential $\phi_{\mu}(\mathrm{T})=\int\psi\circ\mathrm{T}\mathrm{d}\mu$ for $\psi(x)=\frac{1}{2}x^{T}\Lambda^{-1}x$ is equivalent to using a preconditioner with $h^{*}(x)=\frac{1}{2}x^{T}\Lambda x$.
@@
It is well-known that along the Wasserstein gradient flow of the KL divergence starting from a Gaussian and with a Gaussian target (Ornstein-Uhlenbeck process), the measures stay Gaussian [129]. Thus, the Forward-Backward scheme has Gaussian iterates at each step [109, 43]. In this work, we also use a linearly preconditioned Forward-Backward scheme, whose closed-form is derived in (116) (Appendix F). For the Bregman potential, we choose $\phi_{\mu}(\mathrm{T})=\int\psi\circ\mathrm{T}\ \mathrm{d}\mu$ for $\psi(x)=\frac{1}{2}x^{T}\Sigma x$. In this situation, $\mathcal{G}(\mu)=\int V\mathrm{d}\mu$ is 1-smooth and 1-convex relative to $\phi$. Thus, we can apply Proposition 25. We refer to Appendix F for more details on the convexity of $\mathcal{H}$.
@@ G.4 Single-cell experiments
We show the benefits of using the polynomial preconditioner over the single-cell datasets for different metrics.
$\bullet$ The first one considered is the Sliced-Wasserstein distance [102, 16], defined as
@@
$\bullet$ The second one considered is the Sinkhorn divergence [50] defined as
@@
$\mathrm{OT}_{\varepsilon}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\ \int\|x-y\|_{2}^{2}\ \mathrm{d}\gamma(x,y)+\varepsilon\mathrm{KL}(\gamma||\mu\otimes\nu),$ (131)
the entropic regularized OT. The Wasserstein gradient of $\mathrm{S}_{\varepsilon,2}^{2}$ is simply obtained as the potential [50].
$\bullet$ Finally, we also consider the energy distance, defined as
@@ H.1 Proof of Proposition 1
$\quad+o\left(\sqrt{\int\|x-y\|_{2}^{2}\ \mathrm{d}\gamma(x,y)}\right)$
Thus, $\delta\tilde{\mathcal{F}}_{\mu}(\mathrm{S},\mathrm{T}-\mathrm{S})=\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$. Note that in the third equality we used that $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\in L^{2}(\mu)$. ∎
H.2 Proof of Proposition 2
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T}_{\phi_{\mu}}^{\mu,\nu},\mathrm{S})$ $=\phi\big{(}(\mathrm{T}_{\phi_{\mu}}^{\mu,\nu})_{\#}\mu\big{)}-\phi(\mathrm{S}_{\#}\nu)-\int\langle\nabla_{\mathrm{W}_{2}}\phi(\mathrm{S}_{\#}\mu)(y),x-y\rangle\ \mathrm{d}(\mathrm{T}_{\phi_{\mu}}^{\mu,\nu},\mathrm{S})_{\#}\mu(x,y)$
@@
$\geq\mathrm{W}_{\phi}(\nu,\rho)=\mathrm{d}_{\phi_{\rho}}(\mathrm{U}_{\phi_{\rho}}^{\rho,\nu},\mathrm{Id}).$ (135)
In the last line, we used Proposition 15, i.e. that the optimal coupling is of the form $(\mathrm{U}_{\phi_{\rho}}^{\rho,\nu},\mathrm{Id})_{\#}\rho$. ∎
H.3 Proof of Proposition 3
Let $\mathrm{T}_{k+1}=\operatorname*{argmin}_{\mathrm{T}\in L^{2}(\mu_{k})}\ \tau\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{Id})$. Applying the three-point inequality (Lemma 29) with $\psi(\mathrm{T})=\tau\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}$ which is convex, $\mathrm{T}_{0}=\mathrm{Id}$ and $\mathrm{T}^{*}=\mathrm{T}_{k+1}$, we get for all $\mathrm{T}\in L^{2}(\mu_{k})$,
@@ H.3 Proof of Proposition 3
$\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})=\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})-\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{Id})-\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}_{k+1}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}\leq\beta\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})$
~$\iff\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})\leq\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{Id})+\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}_{k+1}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\beta\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id}).$ (138)
@@
~$\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{T}_{k+1})$$\leq\tilde{\mathcal{F}}_{\mu_{k}}(\mathrm{Id})+\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{Id})-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\mathrm{T}_{k+1}).$ (140)
@@
$\mathcal{F}(\mu_{k+1})\leq\mathcal{F}(\mu_{k})-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{Id},\mathrm{T}_{k+1}).\qed$ (142)
H.4 Proof of Proposition 4
@@
$\mathrm{W}_{\phi}(\mu^{*},\mu_{k+1})\leq\left(1-\alpha\tau\right)\mathrm{W}_{\phi}(\mu^{*},\mu_{k})\leq\left(1-\alpha\tau\right)^{k+1}\mathrm{W}_{\phi}(\mu^{*},\mu_{0}).\qed$ (147)
H.5 Proof of Proposition 5
@@
$\phi_{\mu_{k+1}}^{h^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\big{)}$ $=\phi_{\mu_{k}}^{h^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}$
@@
$\quad+\frac{1}{\tau}\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T}_{k+1}),$ (148)
@@
Plugging this into (H.5), we get
@@
$\phi_{\mu_{k+1}}^{h^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k+1})\big{)}\leq\phi_{\mu_{k}}^{h^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}-\frac{1}{\tau}\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id}).\qed$ (153)
H.6 Proof of Proposition 6
Let $\mu^{*}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be the minimizer of $\mathcal{F}$, $k\geq 0$ and $\mathrm{T}=\operatorname*{argmin}_{\mathrm{T}\in L^{2}(\mu_{k}),\mathrm{T}_{\#}\mu_{k}=\mu^{*}}\ \mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T})$. First, observe that since $\mu^{*}$ is the minimizer of $\mathcal{F}$, then $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu^{*})=0$ (see e.g. [74, Theorem 3.1]), and thus $\phi_{\mu_{k}}^{h^{*}}(0)={h^{*}}(0)$. Moreover, it induces that $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{Id},\mathrm{T})=\mathcal{F}(\mu_{k})-\mathcal{F}(\mu^{*})$ and $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{T})=\mathcal{F}(\mu_{k+1})-\mathcal{F}(\mu^{*})$.
@@ H.6 Proof of Proposition 6
[…]
H.7 Proof of Proposition 7
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Since ~$\tilde{\mathcal{F}}_{\mu}^{*}$ is Gâteaux differentiable, we can define its Bregman divergence.
For the first point, $\phi^{h^{*}}$ is $\beta$-smooth relative to $\mathcal{F}_{\mu}^{*}$ along $t\mapsto\big{(}(1-t)\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)+t\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T}\big{)}_{\#}\mu$. Thus, by applying Definition 3 for $s=1$ and $t=0$, we have
$\mathrm{d}_{\phi_{\mu}^{h^{*}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\big{)}\leq\beta\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\big{)}.$ (158)
Using Lemma 19, we finally obtain
$\mathrm{d}_{\phi_{\mu}^{h^{*}}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})\big{)}$ $\leq\beta\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}^{*}}\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{T}_{\#}\mu)\circ\mathrm{T},\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)\big{)}$ (159)
$=\beta\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}\big{(}\mathrm{Id},\mathrm{T}),$
which is the desired inequality.
The second point follows similarly. ∎
H.8 Proof of Lemma 11